20 m/s that makes an angle of 35° with the horizontal. What is the vertical velocity of the projectile after 3 seconds? Take the upward direction to be positive.


1
Expert's answer
2021-01-12T12:10:16-0500

"v_{0y}=v\cdot \sin35°=20\cdot \sin35°=11.47(m/s)"


"v_y(t)=v_{0y}-gt\to v_y(3)=11.47-9.81\cdot 3=-17.96(m/s)" . Answer